# Img-Diff: Contrastive Data Synthesis for Multimodal Large Language Models

### v1

@@ 4. Evaluation on Models trained with Img-Diff
4.1. Overview
To evaluate the effectiveness of our Img-Diff dataset, we use our dataset to finetune the widely used MLLM LLaVA-1.5(Liu et al., 2023b) and MGM(Li et al., 2023d), then evaluate them on extensive benchmarks commonly used for image difference and MLLMs. Specifically, the image difference benchmarks include the MMVP(Tong et al., 2024) benchmark, the Spot-the-Diff(Jhamtani and Berg-Kirkpatrick, 2018) benchmark, and the Image-Edit-Request(Tan et al., 2019) benchmark.
Regarding the finetuning strategy, we mix our data with the original visual instruction tuning dataset of LLaVA-1.5 and MGM respectively and conduct finetuning to get the finetuned MLLMs. For Spot-the-Diff and Image-Edit-Request, since there are training splits in their datasets, we further finetune the MLLMs for an additional 2 epochs using only the benchmark’s training data.
@@ 4.2. Results on the MMVP Benchmark
In Figure 5, we can see that finetuning LLaVA-1.5-7B and MGM-7B with our “object replacement” data noticeably improves their performance on the MMVP benchmark. After fine-tuning, the score of LLaVA-1.5-7B surpass that of LLaVA-1.5-13B. Furthermore, the finetuned MGM-7B show a significant improvement in score compared to the original MGM-7B, even exceeding the scores of the SOTA models GPT-4V and Gemini. This suggests that our dataset enhances the MLLM’s ability to distinguish images with similar CLIP features but different content.
@@ 4.3. Results on the Spot-the-Diff Benchmark
The dataset of Spot-the-Diff comprises a collection of street view images. These images are obtained by capturing scenes from fixed surveillance cameras at different time, resulting in pairs of street view images with minor object differences. Following previous works, our finetuned MLLMs are evaluated on BLEU(Papineni et al., 2002), METEOR(Banerjee and Lavie, 2005), CIDEr-D(Vedantam et al., 2015) and ROUGE-L(Lin, 2004). Table 1 presents the results on the Spot-the-Diff benchmark.
@@ 4.4. Results on Image-Editing-Request
The Image-Editing-Request benchmark is focusing on image editing. Each instance in its dataset consists of an image pair (i.e., a source image and an edited image) and an editing instruction which describes the transformation from the source image to the edited image. During evaluation, our models are required to generate transformation description for these image pairs, and we then calculate the BLEU, METEOR, CIDEr-D, and ROUGE-L scores with the models’ responses and the reference answers. Table 2 presents the results on the Image-Editing-Request benchmark.
@@ 4.5. Results on MLLM Benchmarks
Aside from evaluations related to image difference discrimination capabilities, we also assess the performance of our Img-Diff dataset in enhancing the comprehensive abilities of MLLMs. We test our dataset using commonly used MLLM benchmarks, including VQAv2(Goyal et al., 2017) and GQA(Hudson and Manning, 2019) for assessing the comprehensive VQA capabilities of MLLMs; MMBench(Liu et al., 2023a), MM-Vet(Yu et al., 2023), ScienceQA(Lu et al., 2022), and SEED-Bench(Li et al., 2023c) for testing perceptual and reasoning abilities; and POPE(Li et al., 2023a) for evaluating fine-grained object localization abilities. Table 3 presents the benchmark results on these 8 MLLM benchmarks, with the $\triangle$ metric indicating the percentage improvement averaged across them.
Based on Table 3, it can be observed that after finetuning with our dataset, the performance of LLaVA-1.5-7B shows a comprehensive improvement, with an average increase of 3.06% across all benchmarks. For MGM-7B, the improvements brought by our dataset are not as pronounced as those observed with LLaVA-1.5-7B, but it still achieves an average increase of 1.28%. These score improvements indicate that the finetuned MLLMs not only enhance the ability to discern differences but also improve overall visual capabilities, thereby better addressing VQA tasks.
@@ 4.6. Ablation Studies
To investigate the impact of filtering thresholds on our data performance, we set different filtering thresholds and generate various versions of our dataset. We then finetune multiple versions of LLaVA-1.5-7B using these datasets and evaluate their performance on commonly used MLLM benchmarks. Specifically, the filtering threshold for the Image Similarity Filter of the Difference Area Generator is abbreviated as IS (Image Similarity). The filtering threshold for the Image-Text Matching Filter of the Difference Area Generator is abbreviated as BITM (Bounding Box Image-Text Matching). The filtering threshold for the Caption Similarity Filter of the Difference Captions Generator is abbreviated as CS (Captions Similarity). Lastly, the filtering threshold for the Image-Text Matching Filter of the Difference Captions Generator is abbreviated as CITM (Captions Image-Text Matching). The evaluation results are shown in Table 4.
@@
Based on Table 4, Model (3) adjusts the IS threshold from 0.9-0.98 to 0.85-0.98 compared to Model (4), reducing the filtering intensity for the similarity of image pairs. This adjustment leads to a significant performance decline, indicating that the similarity of image pairs has a substantial impact on data quality. When similarity is low, the data generation process may introduce more noise, as semantic segmentation could generate more areas unrelated to the objects being replaced, resulting in ineffective samples.
@@
Model (2), compared to Model (1), increases the BITM threshold, meaning that when filtering to obtain valid bounding boxes, only those more likely to contain valid objects (i.e., the replaced or replacing objects) are retained. After raising the threshold, slight improvements in model performance are observed. Additionally, Model (5) removes the BITM-related Image-Text Matching filtering compared to Model (3). This operation means that as long as there is any object difference between the areas of the same bounding box of the image pair, the bounding box is considered valid, regardless of whether it contains the valid objects. This operation reduces the filtering intensity, resulting in more instances in our dataset. However, after removing the BITM-related Image-Text Matching filtering, the model’s performance significantly declines. These two experiments demonstrate that higher BITM filtering intensity leads to better model performance, indicating that only bounding boxes related to the replaced or replacing objects should be retained.
@@
Model (3) increases both the CS threshold and the CITM threshold compared to Model (2). Raising the CS threshold implies a greater filtering strength for similar captions, which means that if the two objects inside the two bounding box areas of an image pair have a certain degree of similarity, the bounding box will be filtered out. Increasing the CITM threshold aims to enhance the alignment between the captions and the objects being described. After raising both the CS and CITM thresholds, the model’s performance shows a significant improvement.
From the results shown in Table 4, it can be concluded that the stronger the filtering intensity, the better our dataset’s effectiveness. However, due to the increased filtering intensity resulting in a reduced number of final instances, we choose the settings of Model (3) as our optimal threshold to ensure a sufficient number of generated instances. In our future work, we plan to expand the data sources to generate more pairs of similar images and then evaluate the effects of data obtained with higher filtering intensity.
@@ 5.1. Data Quality
To assess the quality of our Img-Diff dataset, we randomly select 1,000 instances of “object replacement” data and employ multiple professional dataset annotators to evaluate these samples based on three metrics. The final scores are determined through a voting process. Specifically, the first metric is “Bounding Box Difference”, which evaluates whether there are differences between the two sub-images of the same bounding box. If the objects are different, we score it as “high”; if the objects are the same but their features (such as color, shape, etc.) are noticeably different, we score it as “medium”; if the objects are the same and their features are similar, we score it as “low”. The second metric is “Content Caption Accuracy”, which evaluates whether the captions generated by Stage 1 of the Difference Captions Generator accurately describe the sub-images. If both captions are correct, we score it as “high”; if the captions identify the objects but incorrectly describe their features, we score it as “medium”; if the captions incorrectly identify the objects, we score it as “low”. The third metric is “Difference Caption Accuracy”, which evaluates whether the final difference captions accurately describe the object differences between the image pairs. If the description is accurate, we score it as “high”; if the object recognition is correct but the feature description is incorrect, we score it as “medium”; if the object recognition is incorrect, we score it as “low”. The results are shown in Figure 6.
@@ 5.2. Data Diversity
The effectiveness of a dataset is closely related to the diversity of its data. By analyzing the valid object nouns included in the captions of our image pairs, we assess the diversity of our “object replacement” data. Specifically, we count the total number of object categories covered, the total number of unique “object replacement pairs”, and the frequency of occurrences for each object category. The statistical data is presented in Table 5.
Table 5. Diversity of the ”object replacement” data.
@@
As indicated in Table 5, the number of object categories covered by our dataset is 1,203, encompassing most objects in real life. The term “object replacement pair” refers to the combination of the replaced and replacing object names. The total number of unique “object replacement pairs” covered by our dataset is 3,680, with an average of 3.4 occurrences per pair. The total number of occurrences of object categories in our dataset is 25,288, with an average of 21.02 occurrences per category.
To evaluate the coverage of common object categories in our dataset, we analyze the occurrences of the object categories from the Object365(Shao et al., 2019) dataset in our own dataset. The results show that each object category from the Object365 dataset appears an average of 36.07 times in our dataset, totaling 13,164 occurrences, which accounts for approximately 52.06% of the total occurrences of object categories. This indicates that our dataset includes a substantial number of common object categories, ensuring a high frequency of these categories. Additionally, less common object categories make up nearly half of our dataset, demonstrating the broad coverage and high diversity of our dataset.
@@ 6. The “Object Removal” Exploration
Besides object replacement, determining the presence or absence of objects is also crucial. We generate a new set of data focusing on object removal, which prompts MLLMs to analyze which image in a image pair contains the specific object. Finally, we evaluate this new dataset and find that with this data, the performance improvement for LLaVA-1.5-7B increased to 3.91%, while the improvement for MGM-7B reduced to 1.1%. With these results, we conclude that if MLLMs’ finetuning data is limited and does not include data related to object presence, using “object removal” data can be an option for data augmentation. Details of “object removal” data are presented in Appendix A.
@@ 7. Conclusion
In this paper, we draw inspiration from recent advances in contrastive learning and image difference captioning to propose a novel method of contrastive data synthesis, creating a high-quality dataset called “Img-Diff” that focuses on describing object differences. Specifically, we generate many pairs of similar images where the main focus is on object replacement. Then, we use the proposed Difference Area Generator and Difference Captions Generator to generate difference captions for specific regions and form question-answer pairs. In contrast to previous image difference datasets, our dataset focuses exclusively on specific regions inside the images, circumventing the issue where the descriptive information of differences does not fully capture the differences between image pairs, thereby enhancing the accuracy of the data. Afterwards, we fine-tune LLaVA-1.5-7B and MGM-7B using our relatively small-scale dataset, yielding high performance scores on par with SOTA models trained with much larger-scale datasets in image difference tasks, and comprehensive performance improvements in numerous widely recognized MLLM benchmarks. These results confirm that our dataset effectively improves the ability of MLLMs to recognize differences between images and perform detailed image recognition.
@@ A.1. Overview
In the main page, we generate many pairs of similar images focusing on object replacement. Their bounding box regions generally contain objects. However, the ability to determine the object presence is also crucial. Thus, we generate another set of pairs of similar images where the difference lies in the presence or absence of objects, to enhance the model’s ability to determine object presence. We refer to these image pairs as “exist-absent pairs” and the data as “object removal” data.
@@ B.1. Preprocessing of image pairs before inputting into MLLMs
B.2. Model Selection
B.2.1. Overview
@@
B.2.2. Selection of the Semantic Segmentation Model
In our project, we need to use a semantic segmentation model to identify regions in images that may contain objects. To ensure a diverse range of object categories is covered, we opt for models like SAM(Kirillov et al., 2023) instead of traditional semantic segmentation models. Furthermore, to reduce time consumption, we select FastSAM, one of the most efficient and effective models within SAM-like category, as our segmentation model.
B.2.3. Model Size
@@ B.3. Filtering Thresholds
In the Difference Area Generator, after obtaining all effective bounding boxes, we use the IoU method to filter out overlapping bounding boxes. We set the IoU threshold to 0.5, retaining only the bounding boxes with a higher degree of difference for similar positions.
In stage 1 of the Difference Captions Generator, after cropping the images into sub-images and generating content captions, we use the Image-text matching filter to evaluate the matching degree between the sub-images and the content captions. We only consider a caption to be correct if the image-text matching score exceeds 0.4.

### v2

@@ 4. Evaluation on Models trained with Img-Diff
4.1. Evaluation Settings
To evaluate the effectiveness of our Img-Diff dataset, we use it to finetune the SOTA MLLMs, specifically LLaVA-1.5 (Liu et al., 2023b) and MGM (Li et al., 2023d), then evaluate them on extensive benchmarks commonly used for image difference and MLLMs. Specifically, the image difference benchmarks include the MMVP (Tong et al., 2024), the Spot-the-Diff (Jhamtani and Berg-Kirkpatrick, 2018), and the Image-Edit-Request (Tan et al., 2019).
Regarding the finetuning strategy, we mix our data with the original visual instruction tuning dataset of LLaVA-1.5 and MGM respectively and conduct finetuning to get the finetuned MLLMs. For the Spot-the-Diff and Image-Edit-Request benchmarks, since they contain specific training splits, we further fine-tune MLLMs for an additional 2 epochs using only these benchmarks’ training data.
@@ 4.2. Results on the MMVP Benchmark
In Figure 5, we can see that finetuning LLaVA-1.5-7B and MGM-7B with our “object replacement” data noticeably improves their performance on the MMVP benchmark. After finetuning, the score of LLaVA-1.5-7B surpasses that of LLaVA-1.5-13B. Furthermore, the finetuned MGM-7B show a significant improvement in score compared to the original MGM-7B, even exceeding the scores of the SOTA models GPT-4V and Gemini. This suggests that our dataset enhances the MLLM’s ability to distinguish images with similar CLIP features but different content.
@@ 4.3. Results on the Spot-the-Diff Benchmark
The dataset of Spot-the-Diff comprises a collection of street view images. These images are obtained by capturing scenes from fixed surveillance cameras at different times, resulting in pairs of street view images with minor object differences. Following previous works, our finetuned MLLMs are evaluated on BLEU(Papineni et al., 2002), METEOR(Banerjee and Lavie, 2005), CIDEr-D(Vedantam et al., 2015) and ROUGE-L(Lin, 2004). Table 1 presents the results on the Spot-the-Diff benchmark.
@@ 4.4. Results on Image-Editing-Request
The Image-Editing-Request benchmark focuses on image editing. Each instance in its dataset consists of an image pair (i.e., a source image and an edited image) and an editing instruction that describes the transformation from the source image to the edited image. During the evaluation, our models are required to generate transformation descriptions for these image pairs, and we then calculate the BLEU, METEOR, CIDEr-D, and ROUGE-L scores with the models’ responses and the reference answers. Table 2 presents the results of the Image-Editing-Request benchmark.
@@ 4.5. Results on MLLM Benchmarks
Aside from evaluations related to image difference discrimination, we also assess the performance of our “object replacement” data in enhancing the comprehensive abilities of MLLMs. We test our dataset using commonly used MLLM benchmarks, including VQAv2(Goyal et al., 2017) and GQA(Hudson and Manning, 2019) for assessing the comprehensive VQA capabilities; MMBench(Liu et al., 2023a), MM-Vet(Yu et al., 2023), ScienceQA(Lu et al., 2022), and SEED-Bench(Li et al., 2023c) for testing perceptual and reasoning abilities; and POPE(Li et al., 2023a) for evaluating fine-grained object localization abilities. Table 3 presents the results on these MLLM benchmarks, with the $\triangle$ metric indicating the percentage improvement averaged across them.
Based on Table 3, it can be observed that after finetuning with our dataset, the performance of LLaVA-1.5-7B shows a comprehensive improvement, with an average increase of 3.06% across all benchmarks. For MGM-7B, the improvements brought by our dataset are not as pronounced as those observed with LLaVA-1.5-7B, but it still achieves an average increase of 1.28%. These improvements indicate that the finetuned MLLMs not only enhance the ability to discern differences but also improve overall visual capabilities, thereby better addressing VQA tasks.
@@ 4.6. Ablation Studies
To investigate the impact of filtering thresholds on our data performance, we set different filtering thresholds and generate various versions of our “object replacement” dataset. We then finetune multiple versions of LLaVA-1.5-7B using these datasets and evaluate their performance on commonly used MLLM benchmarks. Specifically, the threshold for the Image Similarity Filter of the Difference Area Generator is abbreviated as IS (Image Similarity). The threshold for the Image-Text Matching Filter of the Difference Area Generator is abbreviated as BITM (Bounding Box Image-Text Matching). The threshold for the Caption Similarity Filter of the Difference Captions Generator is abbreviated as CS (Captions Similarity). The threshold for the Image-Text Matching Filter of the Difference Captions Generator is abbreviated as CITM (Captions Image-Text Matching). The evaluation results are shown in Table 4.
@@
Based on Table 4, Model (3) adjusts the IS threshold from 0.9-0.98 to 0.85-0.98 compared to Model (4), reducing the filtering intensity for the similarity of image pairs. This adjustment leads to a significant performance decline, indicating that the similarity of image pairs has a substantial impact on data quality. When the similarity is low, the data generation process may introduce more ineffective instances, as segmentation could generate more areas unrelated to the valid objects (i.e., the replaced or replacing objects).
@@
Model (2), compared to Model (1), increases the BITM threshold, meaning that when filtering to obtain valid bounding boxes, only those more likely to contain valid objects are retained. After raising the threshold, slight improvements in model performance are observed, which demonstrates that only bounding boxes more related to the replaced or replacing objects should be retained.
@@
Model (3) increases both the CS threshold and the CITM threshold compared to Model (2). Raising the CS threshold implies a greater filtering strength for similar captions, which means that if the two objects corresponding to the same bounding box coordinate in an image pair are similar, the bounding box will be filtered out. Additionally, increasing the CITM threshold aims to enhance the alignment between the captions and the objects being described. After raising both the CS and CITM thresholds, the model’s performance shows a significant improvement.
Based on Table 4, it can be concluded that the stronger the filtering intensity, the better our dataset’s effectiveness. However, due to the increased filtering intensity resulting in a reduced number of final instances, we choose the settings of Model (3) as our optimal threshold to ensure a sufficient number of generated instances. In our future work, we plan to expand the data sources to generate more pairs of similar images and then evaluate the effects of data obtained with higher filtering intensity.
@@ 5.1. Data Quality
To assess the quality of our Img-Diff dataset, we randomly select 1,000 instances of “object replacement” data and employ multiple professional dataset annotators to evaluate the samples based on three metrics. The final scores are determined through a voting process. Specifically, the first metric is “Bounding Box Difference”, which evaluates whether there are differences between the two sub-images of the same bounding box coordinate. If the objects are different, we score it as “high”; if the objects are the same but their features (such as color, shape, etc.) are different, we score it as “medium”; if the objects are the same and their features are similar, we score it as “low”. The second metric is “Content Caption Accuracy”, which evaluates whether the captions generated by Stage 1 of the Difference Captions Generator accurately describe the sub-images. If both captions of an image pair are correct, we score it as “high”; if the captions identify the objects but incorrectly describe their features, we score it as “medium”; if the captions incorrectly identify the objects, we score it as “low”. The third metric is “Difference Caption Accuracy”, which evaluates whether the final difference captions accurately describe the object differences between the image pairs. If the description is accurate, we score it as “high”; if the object recognition is correct but the feature description is incorrect, we score it as “medium”; if the object recognition is incorrect, we score it as “low”. The results are shown in Figure 6.
@@ 5.2. Data Diversity
The effectiveness of a dataset is closely related to the diversity of its data. By analyzing the valid object nouns included in the captions of our image pairs, we assess the diversity of our “object replacement” data. Specifically, we count the total number of object categories covered, the total number of unique “object replacement pairs”, and the frequency of occurrences for each object name. The statistical data is presented in Table 5.
Table 5. Diversity of the “object replacement” data.
@@
As indicated in Table 5, the number of object categories covered by our dataset is 1,203, encompassing most objects in real life. The term “object replacement pair” refers to the combination of the replaced and replacing object names. The total number of unique “object replacement pairs” covered by our dataset is 3,680, with an average of 3.4 occurrences per pair. The total number of occurrences for object names in our dataset is 25,288, with an average of 21.02 occurrences per name.
To evaluate the coverage of common object categories in our dataset, we analyze the occurrences of the object names from the Object365(Shao et al., 2019) dataset in our own dataset. The results show that each object name from the Object365 dataset appears an average of 36.07 times in our dataset, totaling 13,164 occurrences, which accounts for approximately 52.06% of the total occurrences of object names. This indicates that our data instances include a substantial number of common object names, ensuring a high frequency of common objects. Additionally, less common object names make up nearly half of our dataset, demonstrating the broad coverage and high diversity of our dataset.
@@ 6. The “Object Removal” Exploration
In addition to object replacement, determining the presence or absence of objects is also crucial. We reform the “object replacement” workflow by integrating an object erasing process, which generates an extended dataset that focuses on object removal. This new data prompts MLLMs to analyze which image in an image pair contains a specific object.
By incorporating this “object removal” data into the “object replacement” dataset, we observed an improvement in performance on eight MLLM benchmarks for LLaVA-1.5-7B, increasing from 3.06% to 3.91%. In contrast, the improvement for MGM-7B slightly decreased from 1.28% to 1.1%. One possible reason for this is that the original training data for MGM-7B has already included a substantial amount of new object existence data compared to LLaVA-1.5.
Based on this exploration, we suggest including “object removal” data as a valuable option to enhance MLLMs, particularly when the finetuning data for MLLMs is limited and lacks sufficient examples related to object presence. More details regarding the ”object removal” data are presented in Appendix A.
@@ 7. Conclusion
In this paper, we draw inspiration from recent advances in contrastive learning and image difference captioning to propose a novel method of contrastive data synthesis, creating a high-quality dataset called “Img-Diff” that focuses on describing object differences. Specifically, we generate many pairs of similar images where the main focus is on object replacement. Then, we use the proposed Difference Area Generator and Difference Captions Generator to generate difference captions for specific regions and form question-answer pairs. In contrast to previous image difference datasets, our dataset focuses exclusively on specific regions inside the images, circumventing the issue where the descriptive information of differences does not fully capture the differences between image pairs, thereby enhancing the accuracy of the data. Afterwards, we finetune LLaVA-1.5-7B and MGM-7B using our relatively small-scale dataset, yielding high-performance scores on par with SOTA models trained with much larger-scale datasets in image difference tasks, and comprehensive performance improvements in numerous widely recognized MLLM benchmarks. These results confirm the effectiveness of our dataset to improve the ability of MLLMs in recognizing detailed differences between images.
@@ A.1. Overview
On the main page, we generate many pairs of similar images focusing on object replacement. Their bounding box regions generally contain objects. However, the ability to determine the object presence is also crucial. Thus, we generate another set of pairs of similar images where the difference lies in the presence or absence of objects, to enhance the model’s ability to determine object presence. We refer to these image pairs as “exist-absent pairs” and the data as “object removal” data.
@@ B.1. Preprocessing of image pairs before inputting into MLLMs
B.2. Training Process for MLLMs
The training process for advanced MLLMs, such as LLaVA-1.5 and MGM, typically involves two stages: pretraining and finetuning. During the pretraining phase, the MLLMs keep the backbone LLM and vision encoder frozen and zero-initialize the learnable projector which is usually an MLP and used for semantic mapping and cross-modality alignment. Only the projector is trained using the pretraining dataset. In the finetuning phase, the MLLMs unfreeze the backbone LLM and finetune both the backbone LLM and the learnable projector using the visual instruction tuning dataset. Specifically, the pretraining dataset is usually an image captioning dataset, while the visual instruction tuning dataset typically consists of VQA datasets for various tasks. Thus, our Img-Diff dataset is integrated into the visual instruction tuning dataset during the finetuning phase and used together with the original dataset to finetune the MLLMs.
B.3. Model Selection
B.3.1. Overview
@@
B.3.2. Selection of the Semantic Segmentation Model
In our project, we need to use a semantic segmentation model to identify regions containing objects in images. To ensure a diverse range of object categories is covered, we opt for models like SAM(Kirillov et al., 2023) instead of traditional semantic segmentation models. Furthermore, to reduce time consumption, we select FastSAM, one of the most efficient and effective models within SAM-like category, as our segmentation model.
B.3.3. Model Size
@@ B.4. Filtering Thresholds
In the Difference Area Generator, after obtaining all effective bounding boxes, we use the IoU method to filter out the overlapping ones. We set the IoU threshold to 0.5, retaining only the bounding boxes with a higher degree of difference for similar positions.
In stage 1 of the Difference Captions Generator, after cropping the images into sub-images and generating content captions, we use the Image-text Matching Filter to evaluate the matching degree between the sub-images and the captions. We only consider a caption to be correct if the image-text matching score exceeds 0.4.